OpenStudy (anonymous):

As light shines from air to another medium, i = 30.0o, the light bends toward the normal and refracts at 22.0o. What is the index of refraction?

OpenStudy (aaronq):

 Snell's law: \(n_1sin\theta_1=n_2sin \theta_2\) \((1)sin(30^o)=n_2sin(22^o)\rightarrow n_2=\dfrac{sin(30^o)}{sin(22^o)}\)